In Compton Scattering, X-rays of wavelength ?i = 0.200 nm are scattered from a block of material. The scattered x-rays are observed at an angle of 45.0° to the incident beam. Calculate the wavelength of the x-rays scattered at this angle.
Answer #1

Soluation Thal Compton Scatte x-vays -The hlave lengt方 of the incide The SCatteédangle.o The Velot The plan냐 Constanr h:G.62
Answer #2

SOLUTION :


Compton shift at an angle 45º :

= Change (increase) in wavelength 

= Compton wave length * ( 1 - cos theta) 

= 0.00243 * 10^(-9) * ( 1 - cos(45))

= 7.117 * 10^(-13) m 


So, wavelength of scattered x-ray 

= 0.200 * 10^(-9) + 7.117 * 10^(-13)

= 0.2007 * 10^(-9) m

= 0.2007 nm (ANSWER)

answered by: Tulsiram Garg